OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x + 3)(x - 1)(x + 5)?

OpenStudy (hockeychick23):

the zeroes are what are added to each x to give you zero, so your zeroes are -3, 1, -5

OpenStudy (campbell_st):

to find the zeros just solve x + 3 = 0 x -1 = 0 x + 5 = 0 you are looking for the values of x, that make the polynomial equal to zero

OpenStudy (anonymous):

Could you help me with this question also? What are the possible rational zeros of f(x) = 3x3 - 4x2 - 7x + 21 ?